### v1

# A Structured Neural ODE Approach for Real Time Evaluation of AC Losses in 3D Superconducting Tapes

Abstract
Efficient modeling of High Temperature Superconductors (HTS) is crucial for real-time quench monitoring; however, full-order electromagnetic simulations remain prohibitively costly due to the strong nonlinearities. Conventional reduced-order methods, such as the Proper Orthogonal Decomposition (POD) and Discrete Empirical Interpolation Method (DEIM), alleviate this cost but are limited by intrusive implementation and by the need for many interpolation points. This work investigates reduced-order strategies for Integral Equation Method (IEM) of HTS systems. We present the first application of PODDEIM to IEM-based HTS models, and introduce a Structured Neural Ordinary Differential Equation (Neural ODE) approach that learns nonlinear dynamics directly in the reduced space. Benchmark results show that the Neural ODE outperforms PODDEIM in both efficiency and accuracy, highlighting its potential for real-time superconducting simulations.
Keywords: High Temperature Superconductor, AC Losses, Integral Equation Method, Proper Orthogonal Decomposition, Discrete Empirical Interpolation Method, Neural ODE
1 Introduction
The second generation (2G) of Rare-Earth Barium Copper Oxide (REBCO)-based HTS has emerged as a promising technology for a wide range of electrical engineering applications. These include power distribution networks, high-field magnets for experimental magnetic confinement fusion devices [1], and particle accelerators [2]. HTS tapes can be configured into several cable architectures, such as the twisted stacked-tape conductor [3], the ROEBEL cable [4], and the Conductor on Round Core (CORC®) cable [5]. Among these, the CORC® design has attracted significant attention due to its advantageous mechanical strength and electromagnetic performance [5].
Although HTS offer the major advantage of operating at significantly higher temperatures than conventional Low Temperature Superconductors (LTS), monitoring the quench phenomenon remains a critical challenge. This difficulty arises from the relatively slow reduction of the critical current $I_{c}$ with increasing temperature [6, 7]. The sudden transition from the superconducting to the normal resistive state generates localized hot spots through the Joule effect, which, if not properly controlled, can lead to irreversible damage of the conductor. Reliable quench detection and mitigation are therefore essential for the safe and continuous operation of future fusion demonstrator power plants, such as the European DEMO [8].
Modeling quench phenomena requires solving multiphysics problems that couple electromagnetics (EM), thermal behavior, and fluid dynamics [9]. While lumped-element models have been proposed, capturing the underlying physics with sufficient accuracy typically demands finite element analysis. For superconducting structures, the EM problem can be addressed using several numerical formulations, most notably the Finite Element Method (FEM) with magnetic field ($H$) formulations [10], or the TTAA approach [11]. An alternative is the IEM, which is traditionally employed in open-domain problems such as antenna radiation, but has also proven effective for computing AC losses in superconductors [12, 13].
Due to the strong nonlinearities in the resistivity model, both FEM and IEM simulations require substantial computational resources to capture the transient, nonlinear dynamics of realistic 3D superconducting structures. This high computational cost renders them impractical for real-time monitoring. To address this limitation, several studies have explored the use of surrogate models to accelerate the EM solution, particularly in the context of AC loss estimation [14, 15]. A first attempt to apply POD to HTS systems within the IEM framework was reported in [16]. However, in this work no hyperreduction strategy was adopted, as the authors argued that the sparsity of the operator $\mathbf{R}$ kept the assembly cost manageable. This is true for relatively small problems, but as the mesh dimension increases, the assembly quickly becomes considerably more demanding. In such cases, the Empirical Interpolation Method (EIM) [17] and its discrete counterpart [18] can be applied to alleviate this cost, and these techniques have already been employed successfully in electromagnetics problems [19]. Nevertheless, the DEIM suffers from two critical drawbacks in the present context: the strong nonlinearity and spatial locality of superconducting material laws require a large number of interpolation points, which reduces efficiency, and its intrusive nature demands full access to the high-fidelity model assembly routines, which limits practical applicability. Furthermore, since the governing equations are highly nonlinear, the reduced-order model must still be solved with an implicit time-integration scheme, which entails a large number of Newton iterations and repeated Jacobian evaluations. As a result, even with DEIM, the simulation cost remains significant.
These challenges have motivated a growing interest in Machine Learning (ML)-based reduction strategies, which are typically non-intrusive and data-driven. Here, we provide only a brief overview of the main approaches, without aiming to be exhaustive; several recent reviews give a comprehensive account of this rapidly evolving area [20, 21, 22].
Neural Operators such as DeepONet [23], the Fourier Neural Operator [24], and graph neural operator variants [25] aim to learn mappings from function spaces, rather than between finite-dimensional vectors. This formulation enables them to generalize across families of input functions, achieving mesh-independent predictions and fast inference once trained. Their flexibility and non-intrusivity make them particularly attractive for parametric problems and large-scale simulations. However, they also present drawbacks: training requires substantial data, the networks can be very large, and their ability to handle complex geometries is limited. Graph-based neural operators alleviate this to some extent, but at the price of predicting the full-order field directly, which implies very large network outputs and nontrivial computational costs.
Physics-Informed Neural Networks (PINNs) [26] incorporate the residual of the physical equation into the training loss, enabling the solution of forward and inverse problems even when data are scarce. PINNs are especially powerful for inverse problems and for settings where measurements are sparse. However, they tend to struggle with complex geometries, and their formulation is not well-suited for input-driven or control problems: a PINN is typically trained for a fixed set of inputs or excitations, and adapting it to new ones generally requires retraining the network.
Operator Inference (OpInf) [27, 28] offers a physics-aware, non-intrusive alternative. Instead of learning black-box surrogates, OpInf identifies reduced operators directly from trajectory data projected onto a reduced basis. This approach has the advantage of being parsimonious, data-efficient, and interpretable, while remaining consistent with the structure of the Reduced Order Model (ROM). Nevertheless, OpInf typically assumes that the reduced operators have a polynomial structure, which can be restrictive in cases where the dynamics involve more complex nonlinear interactions.
Latent Dynamics modeling frameworks [29, 30] take yet another route. The central idea is that the solution trajectories of high-dimensional problems evolve on a low-dimensional manifold, which can be identified either with linear projections (e.g. POD) or with nonlinear encoders (e.g., autoencoders [31, 32]). Once the manifold is fixed, the latent dynamics can be learned in different ways. Sparse Identification of Nonlinear Dynamics (SINDy) [33] discovers sparse governing equations from a library of candidate functions, offering interpretability and parsimony. Neural Ordinary Differential Equations (Neural ODEs) [34] instead parametrize continuous-time dynamics through a trainable vector field.
In the HTS field as well, machine learning–based approaches have recently been applied to accelerate quench detection and AC loss estimation [35, 36]. However, their application to reduced-order modeling of nonlinear electromagnetic simulations remains largely unexplored.
In this work, we build on the latent dynamics paradigm and propose a novel neural architecture for nonlinear electromagnetic modeling that operates directly in the reduced-order space. The starting point is a discretized integral equation model of HTS conductors, which is compressed through POD-based dimensionality reduction. Instead of relying on intrusive hyper-reduction techniques such as DEIM, we design a structured Neural ODE that learns only the nonlinear component of the reduced dynamics, while the linear part is inherited directly from the full-order model. In this way, physical knowledge is embedded into the reduced-order evolution, ensuring parsimony, interpretability, and efficiency. The approach is validated on a HTS CORC® cable case study and benchmarked against the standard PODDEIM method, demonstrating the potential of physics-informed latent dynamics modeling for challenging electromagnetic simulation. Results demonstrate improved accuracy and significantly lower computational cost, highlighting the potential of the proposed strategy for real-time, industrially relevant superconducting applications.
The main contributions of this work are twofold:
  • We demonstrate, for the first time, the application of the POD-DEIM methodology to IEM-based models of HTS systems;
  • We introduce and validate a Structured Neural ODE approach that directly learns the nonlinear dynamics in the reduced-order space, offering a non-intrusive alternative to traditional hyperreduction techniques.
The remainder of this paper is organized as follows. Section 2 describes the physical system under study, including the mesh and material properties. Section 3 introduces the IEM formulation. Section 4 discusses the application of the POD-DEIM method, while Section 5 presents the proposed Neural ODE approach. Section 7 reports the numerical results and Section 6 discusses them. Finally, Section 8 summarizes the conclusions and outlines directions for future research.
2 Model Description
[…]
Figure 1: Resulting 2D mesh of the CORC® cable under analysis.
Within the superconducting domain $\Omega_{sc}\subseteq\Omega$, the resistivity depends on the local current density. In this work, we adopt a power-law constitutive relation [39]:
$\mathbf{E}(\mathbf{J})=\frac{E_{c}}{J_{c}}\bigg(\frac{\lVert\mathbf{J}\rVert}{J_{c}}\bigg)^{n-1}\mathbf{J},$ (1)
where the electric field is assumed to be parallel to the current density. The corresponding resistivity can thus be expressed as
$\rho_{sc}(\mathbf{J})=\frac{E_{c}}{J_{c}}\bigg(\frac{\lVert\mathbf{J}\rVert}{J_{c}}\bigg)^{n-1}.$ (2)
Here $E_{c}=10^{-4}$ V/m is the critical electric field, $J_{c}=236$ MA/m2 is the critical current density, and $n=25$ is the power-law exponent characterizing the superconducting material. In general, both $n$ and $J_{c}$ depend on temperature and magnetic field; here, they are assumed constant for simplicity. Being this constitutive equation highly nonlinear, the resulting problem is also highly nonlinear, which represents a significant challenge when solving it numerically.
We investigate the current density distribution induced by the magnetic field generated by nearby AC cables. Specifically, the magnetic field is assumed to be oriented in the $y$-direction, produced by the current flowing through an AC cable. This external field is modeled as:
$\mathbf{B}_{\text{ext}}(t)=B_{y}(t)\hat{\mathbf{y}}=B_{0}\sin(2\pi ft)\hat{\mathbf{y}}$, (3)
where $B_{0}$ is the magnetic field amplitude, which depends on both the AC current amplitude and the relative distance between the AC cable and the CORC® under study, and $f=50$ Hz is the frequency. According to Faraday’s law of induction, this time-varying magnetic field induces an electric field:
$\nabla\times\mathbf{E}_{\text{ext}}=-\frac{\partial\mathbf{B}}{\partial t}.$ (4)
For the given magnetic field orientation, the induced electric field can be expressed as
$\mathbf{E}_{\text{ext}}(x,y,z)=\begin{bmatrix}-\frac{\partial B_{y}}{\partial t}z\\ 0\\ \frac{\partial B_{y}}{\partial t}x\end{bmatrix}$. (5)
This induced electric field acts as the external excitation driving the current distribution in the superconducting tapes.
3 $J-\varphi_{e}$ Formulation
The $J-\varphi_{e}$ formulation [12] for solving electromagnetic problems is based on the Electric Field Integral Equation (EFIE), which serves as its starting point. We consider an electromagnetic analysis over a bounded domain $\Omega\subset\mathbb{R}^{3}$, where the EFIE is coupled with the current continuity equation:
$\mathbf{E}(\mathbf{r})$ $=-\frac{\partial\mathbf{A}(\mathbf{r})}{\partial t}-\nabla\varphi_{e}(\mathbf{r})+\mathbf{E}_{\text{ext}}(\mathbf{r}),$ (6)
$\ \nabla\cdot\mathbf{J}(\mathbf{r})$ $=0,$
where $\mathbf{E}(\mathbf{r})$ is the electric field, $\mathbf{A}(\mathbf{r})$ is the magnetic vector potential, $\varphi_{e}(\mathbf{r})$ is the scalar electric potential, $\mathbf{J}(\mathbf{r})$ is the current density, and $\mathbf{E}_{\text{ext}}(\mathbf{r})$ is the external applied electric field. The constitutive relation between electric field and current density is given by Ohm’s law:
$\mathbf{E}=\rho(\mathbf{J})\mathbf{J},$ (7)
with $\rho(\mathbf{J})$ denoting the electrical resistivity. The magnetic vector potential $\mathbf{A}$ can be expressed in terms of $\mathbf{J}$ via the integral relation:
$\mathbf{A}(\mathbf{r})=\mu_{0}\int_{\Omega}g(\mathbf{r},\mathbf{r}^{\prime})\mathbf{J}(\mathbf{r}^{\prime})d\Omega,$ (8)
where
$g(\mathbf{r},\mathbf{r}^{\prime})=\frac{1}{4\pi\lVert\mathbf{r}-\mathbf{r}^{\prime}\lVert}$ (9)
is the static Green’s function.
To solve (6) numerically, we adopt the Partial Equivalent Element Circuit (PEEC) methodology. The computational domain $\Omega$ is discretized into triangular elements. The current density and electric potential are expanded using vector basis functions $\mathbf{w}_{h}(\mathbf{r})$ (associated with edges), and piecewise constant basis functions $p_{h}(\mathbf{r})$ (associated with faces):
$\mathbf{J}(\mathbf{r})=\sum_{h=1}^{n_{e}}i_{h}\mathbf{w}_{h}(\mathbf{r}),\quad\varphi_{e}(\mathbf{r})=\sum_{h=1}^{n_{f}}\varphi_{j}p_{h}(\mathbf{r}),$ (10)
where $n_{e}$ is the number of edges and $n_{f}$ the number of faces. Applying Galerkin-testing to (6) yields a system of nonlinear Differential Algebraic Equation (DAE):
$\begin{bmatrix}\mathbf{R}(\mathbf{i})&\mathbf{G}^{T}\\ \mathbf{G}&\mathbf{0}\end{bmatrix}\begin{bmatrix}\mathbf{i}\\ \mathbf{\varphi}\end{bmatrix}+\begin{bmatrix}\mathbf{L}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}\end{bmatrix}\frac{d}{dt}\begin{bmatrix}\mathbf{i}\\ \mathbf{\varphi}\end{bmatrix}=\begin{bmatrix}\mathbf{e}_{s}(t)\\ \mathbf{0}\end{bmatrix}$. (11)
Here $\mathbf{i}\in\mathbb{R}^{n_{e}}$ collects the unknown edge currents, and $\mathbf{\varphi}\in\mathbb{R}^{n_{f}}$ contains the face-associated scalar potentials. The matrix $\mathbf{L}\in\mathbb{R}^{n_{e}\times n_{e}}$ is the inductance matrix, $\mathbf{R}(\mathbf{i}):\mathbb{R}^{n_{e}}\rightarrow\mathbb{R}^{n_{e}\times n_{e}}$ is the current-dependent resistance operator, and $\mathbf{G}$ is the faces-to-edges incidence matrix, acting as discrete counterpart of the divergence operator. The source term $\mathbf{e}_{s}(t)$ arises from the discretization of the external electric field.
Although this formulation leads to non-sparse matrices, it is advantageous because only the regions carrying current—namely, the superconducting domains—need to be meshed. This avoids the need to mesh the surrounding air, unlike the $H$ and $H-\varphi_{e}$ formulations, which require solving for air elements as well. As a result, the overall model size is significantly reduced.
4 POD-DEIM
4.1 Structure-Preserving POD
The POD method [40] is used to derive a reduced-order basis for both current and potential fields while preserving the structure of the underlying equations [41]. The procedure is as follows:
  1. 1.
    Using Full Order Model (FOM) simulations, snapshots of the current and potential are collected:
    $\mathbf{I}=\begin{bmatrix}|&&|\\ \mathbf{i}^{(1)}&\dots&\mathbf{i}^{(n_{s})}\\ |&&|\end{bmatrix}\in\mathbb{R}^{n_{e}\times n_{s}},\quad\mathbf{\Phi}=\begin{bmatrix}|&&|\\ \mathbf{\varphi}^{(1)}&\dots&\mathbf{\varphi}^{(n_{s})}\\ |&&|\end{bmatrix}\in\mathbb{R}^{n_{f}\times n_{s}},$ (12)
    where $n_{e}$, $n_{f}$, and $n_{s}$ are the number of edges, faces, and snapshot instances, respectively.
  2. 2.
    Apply Singular Value Decomposition (SVD) separately to the current and potential snapshot matrices:
    $\mathbf{V}_{i}\mathbf{\Sigma}_{i}\mathbf{U}_{i}^{T}=\mathbf{I},\quad\mathbf{V}_{\varphi}\mathbf{\Sigma}_{\varphi}\mathbf{U}_{\varphi}^{T}=\mathbf{\Phi}.$ (13)
    The columns of $\mathbf{V}_{i}$ and $\mathbf{V}_{\varphi}$ define orthonormal basis vectors for the current and potential, respectively;
  3. 3.
    Retain $r_{i}$ and $r_{\varphi}$ dominant modes, forming the reduced basis
    $\mathbf{V}_{i,r}\in\mathbb{R}^{n_{e}\times r_{i}},\quad\mathbf{V}_{\varphi,r}\in\mathbb{R}^{n_{f}\times r_{\varphi}};$ (14)
    The state variables are projected onto the reduced space as:
    $\mathbf{i}_{r}=\mathbf{V}_{i,r}^{T}\mathbf{i},\quad\mathbf{\varphi}_{r}=\mathbf{V}_{\varphi,r}^{T}\mathbf{\varphi},$ (15)
    and the system matrices are modified accordingly:
    $\mathbf{R}_{r}(\mathbf{i})=\mathbf{V}_{i,r}^{T}\mathbf{R}(\mathbf{i})\mathbf{V}_{i,r},\quad\mathbf{L}_{r}=\mathbf{V}_{i,r}\mathbf{L}\mathbf{V}_{i,r},\quad\mathbf{G}_{r}=\mathbf{V}_{\varphi,r}\mathbf{G}\mathbf{V}_{i,r},\quad\mathbf{e}_{s,r}(t)=\mathbf{V}_{i,r}^{T}\mathbf{e}_{s}(t).$ (16)
This structure-preserving POD approach ensures that the ROM captures the dominant dynamics of both the current and potential while maintaining the original physical structure of the equations.
An a priori estimate of the error introduced by truncating the POD modes is given by the retained energy ratio:
$\epsilon=\frac{\sum_{k=1}^{r}\sigma_{k}}{\sum_{k=1}^{n_{s}}\sigma_{k}},$ (17)
where $\sigma_{k}$ are the singular values of $\mathbf{I}$ or $\mathbf{\Phi}$, stored in descending order in the diagonal of $\mathbf{\Sigma}_{i}$ or $\mathbf{\Sigma}_{\varphi}$, respectively. A higher value of $\epsilon$ indicates a more accurate ROM relative to the full model.
4.2 DEIM
The DEIM [18] efficiently approximates nonlinear functions in ROM s by evaluating them at a small number of carefully selected spatial points. We define the nonlinear function:
$\mathbf{f}(\mathbf{i}):=\mathbf{R}(\mathbf{i})\mathbf{i}.$ (18)
DEIM proceeds by constructing a low-rank approximation of $\mathbf{f}$ through the following steps:
  1. 1.
    snapshots of $\mathbf{f}$ are generated under various input conditions by evaluating (18), using the FOM simulations performed for POD. These snapshots are assembled into a matrix:
    $\mathbf{F}=\begin{bmatrix}|&&|\\ \mathbf{f}^{(1)}&\dots&\mathbf{f}^{(n_{s})}\\ |&&|\end{bmatrix}\in\mathbb{R}^{n_{e}\times n_{s}}.$ (19)
  2. 2.
    as in POD, a SVD is applied to the snapshot matrix:
    $\mathbf{F}=\mathbf{V}_{f}\mathbf{\Sigma}_{f}\mathbf{U}_{f}^{T},$ (20)
    and the reduced basis $\mathbf{V}_{f,r}\in\mathbb{R}^{n_{e}\times r_{deim}}$ is formed from the first $r_{deim}\ll n$ columns of $\mathbf{V}_{f}$;
  3. 3.
    a set of $n_{p}$ optimal interpolation points is determined using a greedy algorithm [17]. These points are encoded in a selection matrix $\mathbf{P}\in\mathbb{R}^{n_{e}\times n_{p}}$.
The number of modes retained, $r_{\text{DEIM}}$, can be chosen based on the retained energy criterion, similar to POD.
Using DEIM, the nonlinear function is approximated as:
$\mathbf{f}(\mathbf{i})\approx\mathbf{V}_{f,r}(\mathbf{P}^{T}\mathbf{V}_{f,r})^{-1}\mathbf{P}^{T}\mathbf{f}(\mathbf{i}).$ (21)
This formulation allows for evaluating the nonlinear term $\mathbf{f}(\mathbf{i})$ only at the selected spatial points determined by $\mathbf{P}$, significantly reducing computational effort.
4.3 POD-DEIM ROM Simulation
We have described how the system is assembled. To simulate its dynamics, we must discretize the system of ODEs in time and solve it step by step. We consider the time step interval $t_{n},t_{n+1}]\subset\mathbb{R}_{\geq 0}$ of length $\Delta t=t_{n+1}-t_{n}$. Given the initial values of the current and potential, $\mathbf{i}_{r}^{(n)}=\mathbf{i}_{r}(t_{n})$ and $\mathbf{\varphi}_{r}^{(n)}=\mathbf{\varphi}_{r}(t_{n})$, our goal is to compute the next-step values $\mathbf{i}_{r}^{(n+1)}=\mathbf{i}_{r}(t_{n+1})$ and $\mathbf{\varphi}_{r}^{(n+1)}=\mathbf{\varphi}_{r}(t_{n+1})$.
$\frac{d\mathbf{i}}{dt}\approx\frac{\mathbf{i}_{r}^{(n+1)}-\mathbf{i}_{r}^{(n)}}{\Delta t},\quad\frac{d\mathbf{\varphi}}{dt}\approx\frac{\mathbf{\varphi}_{r}^{(n+1)}-\mathbf{\varphi}_{r}^{(n)}}{\Delta t}.$ (22)
Using this approximation, we discretize the system with a backward Euler scheme:
$\begin{bmatrix}\Delta t\mathbf{R}_{r}\big(\mathbf{i}_{r}^{(n+1)}\big)+\mathbf{L}_{r}&\Delta t\mathbf{G}_{r}^{T}\\ \Delta t\mathbf{G}_{r}&\mathbf{0}\end{bmatrix}\begin{bmatrix}\mathbf{i}_{r}^{(n+1)}\\ \mathbf{\varphi}_{r}^{(n+1)}\end{bmatrix}=\begin{bmatrix}\mathbf{L}_{r}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}\end{bmatrix}\begin{bmatrix}\mathbf{i}_{r}^{(n)}\\ \mathbf{\varphi}_{r}^{(n)}\end{bmatrix}+\begin{bmatrix}\Delta t\mathbf{e}_{s,r}^{(n+1)}\\ \mathbf{0}\end{bmatrix}$, (23)
with $\mathbf{R}_{r}$ is assembled using only the rows selected by the DEIM procedure. This discretization leads to a nonlinear system, since $\mathbf{R}_{r}$ depends nonlinearly on $\mathbf{i}_{r}^{(n+1)}$. Solving it requires a Newton solver, which in turn involves repeated Jacobian evaluations. This process is computationally expensive. To reduce the cost, we can, in principle, introduce an approximation: we evaluate $\mathbf{R}_{r}$ at the current from the previous time step $\mathbf{i}_{r}^{(n)}$, rather than at the unknown current $\mathbf{i}_{r}^{(n+1)}$:
$\begin{bmatrix}\Delta t\mathbf{R}_{r}\big(\mathbf{i}_{r}^{(n)}\big)+\mathbf{L}_{r}&\Delta t\mathbf{G}_{r}^{T}\\ \Delta t\mathbf{G}_{r}&\mathbf{0}\end{bmatrix}\begin{bmatrix}\mathbf{i}_{r}^{(n+1)}\\ \mathbf{\varphi}_{r}^{(n+1)}\end{bmatrix}=\begin{bmatrix}\mathbf{L}_{r}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}\end{bmatrix}\begin{bmatrix}\mathbf{i}_{r}^{(n)}\\ \mathbf{\varphi}_{r}^{(n)}\end{bmatrix}+\begin{bmatrix}\Delta t\mathbf{e}_{s,r}^{(n+1)}\\ \mathbf{0}\end{bmatrix}$, (24)
where $\mathbf{e}_{s,r}^{(n+1)}=\mathbf{e}_{s,r}(t_{n+1})$, and $\mathbf{R}_{r}$ again assembled from the DEIM-selected rows. This approximation linearizes the system in the unknowns, so advancing the state reduces to solving a linear system—no Newton iterations or Jacobian evaluations are required. However, care must be taken: if the problem is strongly nonlinear, this approximation may introduce large errors, since even a small error in the current can significantly affect $\mathbf{R}_{r}$, and hence the computed next-state current. For the highly nonlinear problem addressed in this paper, this approach proves infeasible, as demonstrated in Section 6.6.
5 Structured Neural ODE
5.1 Model Definition
Instead of approximating the nonlinear operator $\mathbf{R}(\mathbf{i})$ using POD-DEIM, we introduce a neural-network-based surrogate that operates entirely in the POD-reduced current space.
We consider the nonlinear function:
$\mathbf{R}_{r}:\mathbb{R}^{r_{i}}$ $\longrightarrow\mathbb{R}^{r_{i}\times r_{i}}$ (25)
$\mathbf{i}_{r}$ $\longmapsto\mathbf{R}_{r}(\mathbf{i}_{r}),$
that gives the resistance operator in the reduced space as a function of the reduced current-state. This function is modeled as a feed-forward neural network, denoted by $\mathbf{R}_{\psi}$, where $\psi$ represents the set of trainable parameters (weights and biases) of the network. The network consists of sequential layers that apply an affine transformation followed by a smooth, element-wise nonlinearity.
5.2 Training
[…]
$t_{0}$$t_{1}$$t_{2}$$t_{n}$$t_{n+1}$$t_{N-1}$$t_{N}$truestate valuepredictedstate values[$\begin{bmatrix}\mathbf{i}_{r}^{(0)}\\ \mathbf{\varphi}_{r}^{(0)}\end{bmatrix}$$\begin{bmatrix}\mathbf{i}_{r}^{(1)}\\ \mathbf{\varphi}_{r}^{(1)}\end{bmatrix}$$\begin{bmatrix}\mathbf{i}_{r}^{(2)}\\ \mathbf{\varphi}_{r}^{(2)}\end{bmatrix}$$\begin{bmatrix}\mathbf{i}_{r}^{(n)}\\ \mathbf{\varphi}_{r}^{(n)}\end{bmatrix}$$\begin{bmatrix}\mathbf{i}_{r}^{(n+1)}\\ \mathbf{\varphi}_{r}^{(n+1)}\end{bmatrix}$$\begin{bmatrix}\mathbf{i}_{r}^{(N-1)}\\ \mathbf{\varphi}_{r}^{(N-1)}\end{bmatrix}$$\begin{bmatrix}\mathbf{i}_{r}^{(N)}\\ \mathbf{\varphi}_{r}^{(N)}\end{bmatrix}$
Figure 2: Graphical representation of the training process. Green arrows illustrate the discrepancy between the true state values and the predicted values, which the optimization procedure seeks to minimize.
Once the approximated operator $\mathbf{R}_{\psi}$ has been defined, the next step is to train it. A naive approach would be to collect evaluations of $\mathbf{R}$ for different values of $\mathbf{i}$ and then use them to solve a regression problem. However, this strategy is (i) intrusive, since it requires access to snapshots of the nonlinearity, which is often infeasible when dealing with commercial solvers, and (ii) less effective than training on entire trajectories. In fact, as demonstrated in [42], trajectory-based training is superior to local derivative fitting in terms of robustness and generalization capabilities.
For this reason, we adopt a discretize-then-optimize approach, where the time discretization is embedded into the learning process and the parameters are updated by backpropagation through the solver. The procedure to advance the reduced system from $t_{n}$ to $t_{n+1}$ is as follows:
  1. 1.
    Construction of the approximated operator: evaluate the reduced-order operator
    $\mathbf{R}_{\psi}^{(n)}=\mathbf{R}_{\psi}\big(\mathbf{i}_{r}^{(n)}\big),$ (26)
    using the reduced current at the previous time step.
  2. 2.
    Time discretization: apply the Backward Euler scheme to the dynamical system in reduced coordinates.
  3. 3.
    Solution of the linear system: solve the resulting linear system:
    $\begin{bmatrix}\Delta t\mathbf{R}_{\psi}^{(n)}+\mathbf{L}_{r}&\Delta t\mathbf{G}_{r}^{T}\\ \mathbf{G}_{r}&\mathbf{0}\end{bmatrix}\begin{bmatrix}\mathbf{i}_{r}^{(n+1)}\\ \mathbf{\varphi}_{r}^{(n+1)}\end{bmatrix}=\begin{bmatrix}\mathbf{L}_{r}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}\end{bmatrix}\begin{bmatrix}\mathbf{i}_{r}^{(n)}\\ \mathbf{\varphi}_{r}^{(n)}\end{bmatrix}+\begin{bmatrix}\mathbf{e}_{s,r}^{(n+1)}\\ \mathbf{0}\end{bmatrix}$. (27)
    Since $\mathbf{R}_{\psi}$ is evaluated at the previous time step, the resulting system is linear in $\mathbf{i}_{r}^{(n+1)}$ and can be solved without Newton iterations. In this case, considering the previous time current during training allows the network to implicitly account for the resulting lag. Therefore, unlike DEIM, this approximation error is explicitly embedded into the training process: the network learns to compensate for it, thereby reducing the mismatch and improving accuracy.
Repeating this procedure for all time steps $n=0,\dots,N-1$, with $N\Delta t=T$, produced the time evolution of the reduced system. The predicted trajectories $\{\mathbf{i}_{r}^{(n)}\}^{N-1}_{n=0}$, $\{\mathbf{\varphi}_{r}^{(n)}\}^{N-1}_{n=0}$ are then compared with the reference solutions obtained from the FOM, and the parameters $\psi$ are optimized by minimizing the deviation between the reduced and full-order trajectories over the entire time horizon.
This training procedure follows the Neural ODEs [34] paradigm, where the dynamics are unrolled across the temporal domain and the parameters are optimized end-to-end. The loss is defined as the Mean Squared Error (MSE) by the predicted states and the true states obtained by simulating the FOM. Figure 2 provides a graphical interpretation of this training process.
Importantly, inference must employ the same integration scheme used in training, since the learned operator is biased toward that discretization.
6 Results
This section presents the numerical results obtained for the HTS CORC®. We evaluate the performance of both the classical POD-DEIM approach and the proposed Structured Neural ODE model in terms of accuracy and computational efficiency.
6.1 Data Generation
The resulting model features $n_{e}=4518$ edges and $n_{f}=3140$ faces, yielding a total of 7658 Degrees of Freedom (DOFs). To generate the dataset used for POD-DEIM reduction and for training the structured Neural ODE, five transient simulations were performed. Each transient corresponds to a sinusoidal excitation at 50 Hz with different amplitudes. This setup reflects the physical scenario under consideration, where the external magnetic field induced by surrounding cables is sinusoidal at 50 Hz, and its amplitude varies with both current magnitude and cable distance. To evaluate the generalization capability of the proposed approach, additional transient simulations were conducted with varying amplitudes—both within and outside the training distribution—as well as with different excitation frequencies. The corresponding excitation parameters defining the training and testing datasets are reported in Table 1.
Freq. Ampl. Training Within-the-distr. val. Outside-the-distr. val. Different freq. val.
50 Hz 13 mT
50 Hz 15 mT
50 Hz 18 mT
50 Hz 22 mT
50 Hz 24 mT
50 Hz 20 mT
50 Hz 10 mT
50 Hz 30 mT
40 Hz 20 mT
60 Hz 20 mT
100 Hz 20 mT
Table 1: Overview of the excitation parameters adopted for dataset generation. The table reports the combinations of frequencies and amplitudes used in the training simulations, as well as those employed for testing within and outside the training distribution.
Due to the pronounced nonlinearities of the model, time integration was carried out using a fully implicit Backward Euler scheme, with a Newton–Raphson solver applied at each time step. The simulations span a time window from 0 to 35 ms, discretized into 560 steps of duration 0.0625 ms. These runs are computationally expensive: the system is both high-dimensional and nonlinear, requiring many Newton iterations, each involving the solution of a large, dense linear system. All computations were performed on a 13th Gen Intel® Core i7-13700 CPU @2.1 GHz, with 24 cores, equipped with 64 GB of RAM. On this hardware, a single transient required roughly three hours to complete, resulting in about 15 hours for the training set and an additional 18 hours for validation. These times should be understood as indicative rather than absolute, as they depend on the hardware characteristics and implementation details. Importantly, this computational effort belongs to the offline stage of the methodology and must be carried out only once. In practice, the relevant cost is that of repeatedly simulating the ROM, which is several orders of magnitude lower and therefore well suited to design and control tasks.
6.2 POD-DEIM Practical Implementation
In the implementation of the proposed reduction framework, the training snapshots were used for constructing the snapshot matrices. This procedure yielded a total of $n_{s}=2800$ snapshots for the current, potential, and nonlinear term. For the current field, the dominant modes were first selected based on the energy retained, resulting in $r_{i}=48$ retained modes. For the potential field, only the energy retention criterion was applied, leading to the selection of the first five modes, i.e., $r_{\varphi}=5$. The singular values decay for the two variables are shown in Figure 3.
For the DEIM approximation of the nonlinear term, the number of interpolation points $n_{p}$ was chosen to be equal to the number of retained modes. In this case, 150 modes were retained. The relatively large number of DEIM modes is motivated by the strong nonlinearities present in the system: accurately capturing localized variations in the current density requires a correspondingly dense set of interpolation points. The singular values decay for the nonlinear term, with the details of retained modes, are shown in Figure 3.
Figure 3: Singular value decay of the current, potential, and nonlinear term $\mathbf{R}(\mathbf{i})\mathbf{i}$ snapshot matrices. Retained modes are shown in orange, discarded modes in blue.
6.3 Structured Neural ODE Practical Implementation
The neural network $\mathbf{R}_{\psi}$ is implemented as a fully connected feed-forward architecture with $L_{h}=4$ hidden layers, each consisting of 140 neurons. As a preprocessing step, the absolute value of the reduced current $|\mathbf{i}_{r}|$ is fed to the network, reflecting the fact that resistivity—and therefore the $\mathbf{R}_{\psi}$—depends on the magnitude of the current rather than its sign. All the layers employ selu [43] activation functions. The training dataset considers the very same transients used for constructing the POD-DEIM snapshot matrix. This dataset is split into fixed-length sequences that are organized into mini-batches, and the model is trained using the Adam optimizer [44].
6.4 Accuracy
The accuracy of each reduced-order model was assessed by comparing the predicted current density with the reference solution obtained from the FOM. Specifically, the instantaneous absolute error was computed as the absolute difference between the ROM-predicted current and the corresponding FOM value. In the time-dependent plots, the mean, 95th-percentile, and maximum errors were evaluated across all current states at each time instant, thereby illustrating the temporal evolution of the error. Conversely, in the summary tables, the reported mean, 95th-percentile, and maximum values were computed over all current states and all time instants of the transient simulations.
6.4.1 Within-the-distribution Validation
To assess the accuracy of the proposed approach within the training distribution, the models were evaluated using sinusoidal inputs whose amplitude and frequency fall inside the range of the training data ($f=50$ Hz, $B_{0}=20$ mT). The POD-DEIM ROM was simulated with the same discretization strategy employed for the FOM, whereas the structured Neural ODE ROM was advanced according to (27). As illustrated qualitatively at the top of Figure 4, both reduced-order models reproduce the state evolution with good accuracy.
A quantitative comparison with the FOM simulation further supports this result: the bottom part of Figure 4 shows the mean, 95th-percentile, and maximum error among the states as functions of time, while the corresponding numerical values are summarized in Table 2. These results indicate that the errors are of the same order of magnitude, with the structured Neural ODE yielding slightly lower values.
Figure 4: Top: Evolution of selected current states in the within-the-distribution validation transient for the ROM simulations (dashed red) compared to the FOM (solid blue). Bottom: Corresponding mean, 95th-percentile, and maximum errors between the ROM s and the FOM.
DEIM Neural ODE
Mean Error 0.1709 A/m 0.1164 A/m
95th-percentile 0.7921 A/m 0.3398 A/m
Max Error 4.5680 A/m 1.7775 A/m
Table 2: Mean, 95-th percentile and max absolute error for the DEIM ROM and the Neural ODE ROM in the within-the-distribution validation transient.
Moreover, the current density map obtained with the proposed Neural ODE approach is consistent with that from the FOM simulation, as shown in Figure 5, which depicts the current distribution and the associated error at selected time instants.
Figure 5: Comparison between the current density map of the FOM simulation and the Neural ODE-based ROM on the within-the-distribution validation transient. The results demonstrate high accuracy.
6.4.2 Outside-the-distribution Validation
To evaluate the performance of the ROM s outside the training distribution, two additional transient simulations were performed at 50 Hz with excitation amplitudes of 10 mT and 30 mT—values lying below and above the training range (13–24 mT), respectively. The corresponding quantitative results, summarized in Table 3, indicate that the accuracy of the structured Neural ODE decreases noticeably, whereas the POD-DEIM ROM maintains performance comparable to the within-distribution case.
DEIM Neural ODE
Mean Error 0.1021 A/m 0.4566 A/m
95th-percentile 0.3860 A/m 2.0990 A/m
Max Error 6.0196 A/m 6.7252 A/m
Table 3: Mean, 95-th percentile and max absolute error for the DEIM ROM and the Neural ODE ROM in the outside-the-distribution validation transients.
This outcome is consistent with expectations: data-driven models such as the structured Neural ODE are inherently constrained by the range of the training data. When the excitation amplitude extends beyond this range, the resulting state trajectories explore regions of the state space not represented during training, leading to a loss in predictive accuracy.
6.4.3 Different Frequency Validation
To further assess the generalization capability of the proposed approach, additional simulations were carried out at an excitation amplitude of 20 mT—within the training range—but with frequencies of 40, 60, and 100 Hz, differing from the 50 Hz training frequency. The quantitative results, summarized in Table 4, demonstrate that the structured Neural ODE achieves very good accuracy across all tested frequencies, with errors of the same order of magnitude as in the 50 Hz validation and consistently lower than those of the POD-DEIM ROM.
DEIM Neural ODE
Mean Error 0.1729 A/m 0.1484 A/m
95th-percentile 0.7854 A/m 0.4595 A/m
Max Error 4.6427 A/m 2.7591 A/m
Table 4: Mean, 95-th percentile and max absolute error for the DEIM ROM and the Neural ODE ROM in the different frequency validation transients.
This strong performance can be attributed to the intrinsic structure of the proposed model: since the Neural ODE component depends solely on the state variables, the learned dynamics remain accurate as long as the system evolution explores regions of the state space already represented in the training data. Although the model implicitly captures some temporal dependencies through the state update formulation, these do not significantly affect the accuracy, explaining the robust behavior observed across different excitation frequencies.
6.5 Computational Cost
To compare the computational cost of the two ROM s, we evaluate the number of Floating-Point Operations (FLOPs) required to advance the system by one time step. The time-stepping scheme used to simulate the POD-DEIM ROM employs Newton iterations at each step, with residual and Jacobian evaluations at each iteration—an approach necessitated by the strong nonlinearity of the problem. In both cases, an optimal implementation scenario is assumed.
6.5.1 POD-DEIM ROM Computational Cost
The computational cost of the POD-DEIM reduced-order model involves mapping the reduced solution to the full current density $\mathbf{J}$, evaluating the resistivity $\rho$ and its derivative $\rho^{\prime}$, assembling the DEIM-selected rows of $\mathbf{R}$ and its Jacobian, performing matrix-vector products and projections, and solving the linear system within each Newton-Raphson iteration. Contributions from the right-hand side, which are independent of the solution, are computed once per time step.
6.5.2 Structure Neural ODE ROM Computational Cost
The computational cost for advancing the state in the POD-DEIM reduced-order model includes evaluating the neural network, assembling the left- and right-hand sides through matrix-vector operations, and solving the resulting linear system.
6.5.3 Results of the Evaluations
For a single time step, the POD-DEIM ROM requires approximately $4.5\times 10^{7}$ operations, while the structured Neural ODE ROM requires only $2.6\times 10^{5}$, corresponding to a theoretical speed-up of $\times 176$. However, achieving this theoretical performance depends on highly optimized implementations. While optimizing the Neural ODE implementation is relatively straightforward—requiring only efficient neural network evaluation and basic algebraic operations—achieving an optimal DEIM implementation is more challenging. In fact, when using DEIM, the FOM assembly routines are still called on the fly during the reduced-order simulation. To reach optimal efficiency, these routines themselves would need to be carefully optimized, which is a nontrivial task in practice. In our implementation, the practical speed-up reaches $\times 1500$, reducing computation from 5 minutes to just 0.2 seconds for a full transient reconstruction. These results are summarized in Table 5.
Approach # FLOPs Eval. Time
DEIM 45,330,701 ($\times 176$) 334 s ($\times 1500$)
Neural ODE 257,395 ($\times 1$) 0.2 s ($\times 1$)
Table 5: Number of FLOPs per time step and total evaluation time for reconstructing a full transient.
6.6 Note on the Time-stepping Strategy
As explained in Section 4.3, the DEIM approach could also be implemented by evaluating $\mathbf{R}_{r}$ at the previously computed state, in which case only a linear system would need to be solved at each time step. This would reduce the computational effort to a level comparable with that of our Neural ODE. In particular, if we consider this approximation, the DEIM requires approximately $1.99\times 10^{5}$ FLOPs to advance the state, which is about 1.29 times fewer operations than the structured Neural ODE. However, a crucial difference lies in the fact that the neural ODE has been explicitly trained to perform this task, and the training process inherently accounts for this simplification, thereby ensuring accurate results. By contrast, due to the strong nonlinearity of the problem, applying the same type of approximation within the POD-DEIM ROM leads to unreliable and ultimately uninformative simulations, as illustrated in Figure 6.
Figure 6: Evolution of selected current states in the validation transient for the POD-DEIM ROM (dashed red) compared with the FOM (solid blue). The corresponding mean, 95th-percentile, and maximum errors between the ROM and the FOM are also reported. In these simulations, the nonlinear operator is evaluated using the current state from the previous iteration. The results clearly show that this approach does not yield accurate or practically useful solutions.
6.7 Post-processing: Power losses
As a post-processing step, we evaluate the AC losses in the CORC® cable, which serve as input for a thermal model aimed at predicting the cable’s temperature distribution—a key quantity for quench prediction in both design and control. The AC losses are computed as:
$p(t)=\int_{\Omega_{sc}}\mathbf{J}\cdot\mathbf{E}d\Omega=\int_{\Omega_{sc}}\rho\lVert\mathbf{J}\rVert^{2}d\Omega.$ (28)
These losses are related to the current loops in the HTS tapes, also recognized as magnetization losses [45]. The integral is performed numerically, assuming the electric field and current density are constant within each element. Power losses are evaluated using current densities obtained from both the POD-DEIM ROM and the Neural ODE ROM. Figure 7 shows the results of these evaluations, along with the error relative to the FOM values. The comparison indicates that both ROM s achieve comparable accuracy in predicting power loss.
Figure 7: Comparison of power loss of the within-the-distribution validation transient computed from the current density obtained using the FOM and the two ROM s, including the associated error.
7 Discussion
The comparison between the structured Neural ODE and the POD-DEIM approaches highlights a clear trade-off between accuracy, computational efficiency, intrusiveness, and training requirements.
Both ROM s achieve comparable overall accuracy, with the structured Neural ODE generally providing slightly higher precision. In particular, the structured Neural ODE yields the most accurate results for inputs within the training distribution, while the POD-DEIM ROM proves more reliable when the excitation amplitude lies outside the training range. For excitations at different frequencies, the structured Neural ODE performs remarkably well, maintaining accuracy comparable to the training frequency case and outperforming the POD-DEIM model. This behavior confirms that the proposed data-driven structure generalizes effectively across variations in frequency, provided that the system trajectories remain within regions of the state space represented during training.
In terms of computational cost, the structured Neural ODE is significantly faster because it advances the state by solving only a linear system at each time step, whereas the POD-DEIM ROM requires a high number of Newton-Raphson iterations due to the strong nonlinearity of the model, with repeated evaluations of the Jacobian at each step.
The structured Neural ODE is also considerably less intrusive, as it only requires access to the assembled linear operators and to evaluations of the nonlinear operator at selected working points. Such information can typically be extracted without modifying the underlying simulation code, making the approach compatible with commercial solvers and straightforward to integrate into existing workflows. In contrast, the POD-DEIM method is highly intrusive, since it requires direct access to and modification of the FOM assembly routines to incorporate the DEIM interpolation points. This need for code-level intervention limits its applicability in commercial or legacy software environments and makes the optimization of the DEIM approximation more demanding in practice.
When it comes to training, POD-DEIM is easier to set up and tune due to its linear algebra-based structure, whereas structured Neural ODEs demand careful tuning of hyperparameters and more complex training procedures.
Finally, data requirements differ: POD-DEIM typically needs less data and generalizes well across operating conditions, while structured Neural ODEs require larger, representative datasets to ensure reliable performance during operation. A concise summary of these pros and cons is provided in Table 6.
Overall, these considerations suggest that while POD-DEIM may be advantageous in simpler or data-scarce scenarios, the structured Neural ODE offers superior speed, flexibility, and precision in settings where sufficient data are available and intrusiveness must be minimized.
Aspect Neural ODE ROM POD-DEIM ROM
Accuracy High (overall, slightly higher) High
Cost Low ($\sim$176× speedup) High (Newton Iterations; Jacobian Evaluations)
Data requirements High (Dense Dataset) Low (Few snapshots are enough)
Training procedure Empirical and delicate Straightforward (Linear Algebra-based)
Intrusiveness Non-intrusive Intrusive
Table 6: Detailed comparison between Neural ODE-based ROM and POD-DEIM ROM.
8 Conclusion
In this work, we developed and assessed reduced-order modeling strategies for nonlinear EM simulations of HTS conductors within the IEM framework. Specifically, we presented the first application of POD-DEIM to IEM-based HTS models, and proposed a Structured Neural ODE approach that learns the nonlinear dynamics directly in the reduced-order space.
These methods were pursued to enable real-time prediction and monitoring of quench phenomena in superconducting devices. While conventional FEM or IEM remain too costly for online use, the proposed reduced-order strategies significantly improve computational efficiency without sacrificing accuracy. In particular, the Structured Neural ODE demonstrated superior predictive capability compared to the standard POD-DEIM approach, achieving faster evaluations while maintaining high fidelity to the FOM.
Future work will extend this framework toward multiphysics simulations, where the computed AC losses from the electromagnetic model will serve as inputs to thermal solvers. This coupling will allow for a more comprehensive description of quench initiation and propagation, thereby bringing reduced-order modeling closer to practical deployment in the design and operation of next-generation superconducting systems.